OpenStudy (anonymous):

square root3x-7 = square root 2x+5 solve the equation???

OpenStudy (anonymous):

\[\sqrt{3x-7} = \sqrt{2x+5}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the sqrt cancel out

OpenStudy (anonymous):

\[ 3x-7 = 2x+5\]

OpenStudy (anonymous):

you should be able to do it now

OpenStudy (anonymous):

\[ 3x-2x = 5+7\]

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

awesome thank you